# Common Envelope Evolution of Ultralight Boson Clouds

### v1

@@ V.1 Initial Conditions
Superradiant clouds typically grow individually around their own host black holes when the orbit is in a very large separation. In this case, each cloud is mainly in one of its atomic states; therefore, the initial wavefunction is naturally given by the superposition of the two atomic states. To utilize the framework built in Sec. III, we need to decompose such an initial state into molecular eigenstates. To be concrete, we consider two identical black holes inspiral in circular orbit with their spins aligned with the orbital angular momentum. The clouds associated with each black hole are assumed to be dominated by a same atomic eigenstate, and hence the initial wavefunction of the two clouds can be written as
@@
where $E_{\rm nlm}$ is the energy of the atomic eigenstate, and $\varphi_{a,\rm nlm}$ is the atomic eigenstate centered on the $a$th-black hole. Note that $\varphi_{a,\rm nlm}$ grew via superradiance, and is different from the eigenstate $\bar{\varphi}_{n_{s}km}$ in Eq. (12). Here we use non-italic quantum numbers to indicate the difference. In particular, $\rm m$} characterizes the angular momentum of $\varphi_{a,\rm nlm}$ along the black hole spin, while $m$ characterizes the angular momentum along $\mathbf{r}_{2}-\mathbf{r}_{1}$. To relate the initial state $\Psi_{0}$ with the co-rotating molecular eigenstates $\bar{\Psi}_{n\ell m}$, we first decompose $\varphi_{a,\rm nlm}$ into angular momentum eigenfunctions in the direction of $\mathbf{r}_{2}-\mathbf{r}_{1}$, i.e., $\bar{\varphi}_{a,\rm nlm}$. Doing so introduces a time-dependent phase and also mixing between different angular momentum eigenstates,
@@
where $c_{n\ell m}^{0}$ are constants and the sum of $\bar{\Psi}_{n\ell m}$ only involves a subset of molecular eigenstates. Moreover, in the large separation limit, we have $E_{\rm nlm}\approx\epsilon_{i}$ (cf. Fig. 2). Thus, we conclude that the initial state can be written as a sum of several co-rotating molecular eigenstates with a time dependent phase,
@@
For example, if the initial cloud consist of two spin-aligned states $\varphi_{211}$, we have
@@
where the eigenstate wavefunctions are normalized to 1. Here we discuss only the case with two identical gravitational atoms. If the clouds surrounding the two black holes are in different atomic eigenstates or with non-parallel spin alignment, the initial condition can also be constructed in a similar way for equal mass binaries. Going beyond equal mass binaries is, however, non-trivial. The initial conditions for $c_{i}$ in Eq. (14) cannot be obtained analytically. Nevertheless, it can be obtained by projecting the initial wavefunction on the molecular eigenstates numerically. The initial conditions for clouds in unequal mass binaries will be left for future investigation.
@@ V.2 Cloud Depletion
Before forming a common envelope, clouds in binaries may undergo mass transfer. For example, cloud that is initially in $\varphi_{1,{\rm 211}}$ starts to oscillate between $\varphi_{1,{\rm 211}}$ and $\varphi_{2,{\rm 21-1}}$ when the wavefunctions of the two states develop a noticeable overlap as the two black holes get closer. As a result, the cloud loses mass to the second black hole via $\varphi_{2,{\rm 21-1}}$. It is claimed in Ref. [44] that, for equal mass binaries, the cloud depletes rapidly around $a\sim 60r_{B}$ due to mass transfer to the $\varphi_{{\rm 21-1}}$ mode of the companion black hole. In this case, the cloud cannot survive to form a common envelope. In the following, we will show that the depletion is not efficient, except for very large $q$.
@@
During non-resonant evolution, we have
@@
where $\eta$ is the coupling strength, $\omega\approx\Delta\epsilon+2\Omega$ with $\Delta\epsilon$ being the energy split between $\varphi_{1,{\rm 211}}$ and $\varphi_{2,{\rm 21-1}}$, and the integral over $t^{\prime}$ takes into account the slowly changing of $\omega$ and $\eta$. For $q\neq 1$, we have $|\Delta\epsilon|\approx(1-q^{2})\alpha^{3}/8\gg|2\Omega|$, and the cloud deplete significantly only if
$\Gamma_{2,{\rm 21-1}}\frac{\eta^{2}}{\omega^{2}}\frac{\Omega}{\dot{\Omega}}\approx\frac{40}{3}\frac{(1+q)^{1/3}q^{8}}{(1-q^{2})^{2}}\alpha^{2}{\rm I}_{4}(a)\gtrsim 1.\,$ (37)
where we have averaged $|c_{2}|^{2}$ over a time scale much longer than $\sqrt{\omega^{2}+\eta^{2}}$ but shorter than $\Omega/\dot{\Omega}$. We also defined ${\rm I}_{4}(a)=\left(\frac{a}{r_{B}}\right)^{4}G^{2}M^{2}\eta^{2}/\alpha^{6}$ which depends only on $q$. Fig. 5 shows the numerical value of ${\rm I}_{4}(a)$ for different $q$, from which we conclude that the cloud barely deplete unless $q$ is relatively large. For example, by requiring the maximum of l.h.s. of the inequality (37) is less than $1$, we get $q>3.27$ for $\alpha=0.1$. The case of $q=1$ is specially, in the sense that $|\Delta\epsilon|\simeq 0$, and $\omega$ is dominated by $2\Omega$. [^4] To be precise, $|\Delta\epsilon|\propto\mu\alpha^{5}$ and depends on the black hole spin. In this case, the depletion condition becomes
@@
where ${\rm I}_{7}(a)=\left(\frac{a}{r_{B}}\right)^{7}G^{2}M^{2}\eta^{2}/\alpha^{6}$. As shown in Fig. 5, even for $q=1$, the cloud cannot completely deplete. Compared to the estimation in Ref. [44], $\omega$ differs by $2\Omega$, which is due to binary rotation that is not considered in Ref. [44]. Moreover, the mass transfer can be resonant at certain orbits as discussed in Ref. [29], which, however, do not deplete the cloud completely either. Therefore, we conclude that, for the general mass ratio, the clouds generically survive to form common envelopes in the late inspiral.
Figure 5: Integrals in Eq. (37) and Eq. (38). The vertical axis shows ${\rm I}_{x}(a)=\left(\frac{a}{r_{B}}\right)^{x}G^{2}M^{2}\eta^{2}/\alpha^{6}$, where colors of curves represent conditions for various mass ratios ($q=0.5$, $0.75$, 11, $1.25$ and $1.5$ ) with corresponding powers.
@@ VI Implications on orbital evolution
$E_{\rm c}=M_{\rm c}\alpha^{2}\left[\sum_{i}|c_{i}|^{2}\tilde{\epsilon_{i}}+(1+q)^{1/2}\tilde{a}^{-3/2}\sum_{i,j}c_{i}^{*}c_{j}\tilde{\eta}_{ij}\right]$ (39)
where $M_{c}$ is the cloud mass, and $\tilde{\epsilon}_{i}$ and $\tilde{\eta}$ are the dimensionless molecular eigenenergy and coupling strength, for example, shown in Fig. 2 and Fig. 3 respectively. The cloud also contributes to the total angular momentum with
@@
Figure 6: Orbital evolution in the phase space. The upper plot shows the evolution trajectories in the absence of clouds, where the eccentricity always decreases during the orbital decay. The lower plot shows the evolution trajectories in the presence of the cloud, where we assume a cloud of $M_{c}=0.1M$, originating from two spin-aligned superradiance states $\varphi_{1,211}$ and $\varphi_{2,211}$, and undergoing adiabatic evolution. In this case, the eccentricity may increase as the orbits decay.
@@
According to Eq. (42), the cloud plays an important role in the eccentricity evolution during orbital decay. In particular, the rhs of Eq. (42) can be positive in the presence of the cloud, resulting in increased orbital eccentricity. For example, we consider the cloud that initially consists of two spin-aligned atomic states $\varphi_{1,211}$ and $\varphi_{2,211}$. During adiabatic evolution, we expect that the coefficients $c_{i}$ are constant in time and that the orbital evolution can be obtained by solving Eq. (41) and Eq. (42). In Fig. 6, we plot the evolution trajectories, assuming $\tilde{M}_{c}=0.1$. We find that eccentricity may increase when $4<\tilde{a}<6$ and $10<\tilde{a}$.
Figure 7: Eccentricity pumping during molecular level transitions. The plots show orbital evolution near the molecular level transition from $\bar{\psi}_{211}$ to $\bar{\psi}_{210}$, assuming different cloud masses. The upper plot assumes an initial eccentricity of $0.2$ at $\tilde{a}=3$, while the lower plot assumes an initial eccentricity of $0.1$ at $\tilde{a}=3$. The gray dashed line located at $a\sim 2.79r_{B}$ represents the resonant radius of the transition. The evolution is obtained by solving Eq. (41), Eq. (42), and Eq. (43) numerically.
The growing eccentricity may also occur during nonadiabatic evolution, where $c_{i}$ change significantly due to resonant transitions of the cloud. For the cloud considered above, there is a resonant transition from $\bar{\psi}_{211}$ to $\bar{\psi}_{210}$ at $\tilde{a}\simeq 2.79$. In Fig. 7, we show the eccentricity evolution near the resonant orbit by solving Eq. (41), Eq. (42) and Eq. (43) numerically. We find that for $\tilde{M}_{c}>10^{-2}$, the eccentricity vary significantly during the transition, leaving an eccentricity of ${\cal O}(0.1)$ at $a\sim r_{B}$.
Assuming equal mass binaries, the frequency of the gravitational wave emitted at $a\sim r_{B}$ can be estimated as
@@
Therefore, for stellar mass binaries, the superradiant cloud can give rise to a non-zero eccentricity detectable in the LVK band. Interestingly, recent analysis on the LVK event GW200105 indicates a median orbital eccentricity of $e\sim 0.145$ at an orbital period of $0.1$s [52]. Our work provides a potential origin of the median eccentricity. [^5] Note that, in practice, the mass of the boson field is a constant, and $\alpha$ varies with black hole masses. In other words, one has $f_{B}\propto\mu^{3}M^{2}$. Nevertheless, we do not expect to have eccentricity excitation in binaries with heavier black holes, because the superradiant growth is suppressed or the cloud might have already depleted. Therefore, the eccentricity excitation mechanism could explain the observed eccentricity in GW200105, while being in consistence with other LVK events.
Note that the eccentricity excitation mechanisms discussed above are different from the one in Refs. [42, 43], which is caused by the level transitions between atomic states, and happens only in inclined orbit and typically at relatively large orbital separation. As a result, the induced eccentricity hardly survives to the LVK band, and cannot be detected with the current ground based gravitational wave detectors.
@@ VII Conclusion and Discussion
In this work, we investigate the evolution of the cloud in the stage where the orbital separation becomes comparable to the cloud size. We describe the cloud evolution with the gravitational molecular eigenstates. With this framework, we discuss the common envelope formation of the superradiant clouds and their late evolution in binaries. We find that most of the time the cloud evolves adiabatically, i.e. each eigenstate slowly adjusts itself as the orbit decays. Resonant transitions between different eigenstates may also happen when the energy levels cross or if the orbital frequency matches the energy difference between the eigenstates.
We have also investigated the potential effects on the orbital evolution of the clouds. We find that for elliptical orbits, the eccentricity may increase when the orbital separation is greater than 5 times the Bohr radius. This result proposes a potential origin for the median eccentricity at small orbital separation that was recently implied by the LVK event GW200105 [52]. The cloud induced eccentricity could be further tested with future gravitational-wave observation, providing an alternative way of searching for the ultralight boson field with gravitational wave observations.
This work provides a framework for investigating cloud evolution and its effects on binary black hole mergers as environments. Although in this work, we focus on the conservative dynamics, the bound states could also couple with the unbounded states by $H^{\prime}$ and excite radiation of the boson field, leading to cloud ionization and extra dissipation of the binary binding energy. Ionization of cloud in comparable mass black hole binaries has been recently observed in numerical simulations and analyzed in the recent work [53]. Our work provides an alternative way of investigating the ionization of the gravitational molecule. Together with the ionization process, the cloud environmental effects on gravitational waves from binary black hole mergers are promising directions for future study.
@@ ACKNOWLEDGMENTS
The authors thank Yifan Chen and Giovanni Maria Tomaselli for useful discussions. J. Z. is supported by the scientific research starting grants from the University of Chinese Academy of Sciences (Grant No. 118900M061), the Fundamental Research Funds for the Central Universities (Grants No. E2EG6602X2 and No. E2ET0209X2), and the National Natural Science Foundation of China (NSFC) under Grant No. 12147103.

### v2

@@ V.1 Initial conditions
Superradiant clouds typically grow individually around their own host black holes when the orbit is in a very large separation. In this case, each cloud is mainly in one of its atomic states; therefore, the initial wave function is naturally given by the superposition of the two atomic states. To utilize the framework built in Sec. III, we need to decompose such an initial state into molecular eigenstates. To be concrete, we consider two identical black holes inspiral in circular orbit with their spins aligned with the orbital angular momentum. The clouds associated with each black hole are assumed to be dominated by the same atomic eigenstate, and hence the initial wave function of the two clouds can be written as
@@
where $E_{\rm nlm}$ is the energy of the atomic eigenstate, and $\varphi_{a,\rm nlm}$ is the atomic eigenstate centered on the $a$th black hole. Note that $\varphi_{a,\rm nlm}$ grew via superradiance, and is different from the eigenstate $\bar{\varphi}_{n_{s}km}$ in Eq. (12). Here we use nonitalic quantum numbers to indicate the difference. In particular, $\rm m$} characterizes the angular momentum of $\varphi_{a,\rm nlm}$ along the black hole spin, while $m$ characterizes the angular momentum along $\mathbf{r}_{2}-\mathbf{r}_{1}$. To relate the initial state $\Psi_{0}$ with the corotating molecular eigenstates $\bar{\Psi}_{n\ell m}$, we first decompose $\varphi_{a,\rm nlm}$ into angular momentum eigenfunctions in the direction of $\mathbf{r}_{2}-\mathbf{r}_{1}$, i.e., $\bar{\varphi}_{a,\rm nlm}$. Doing so introduces a time-dependent phase and also mixing between different angular momentum eigenstates,
@@
where $c_{n\ell m}^{0}$ are constants and the sum of $\bar{\Psi}_{n\ell m}$ only involves a subset of molecular eigenstates. Moreover, in the large separation limit, we have $E_{\rm nlm}\approx\epsilon_{i}$ (cf. Fig. 2). Thus, we conclude that the initial state can be written as a sum of several corotating molecular eigenstates with a time-dependent phase,
@@
For example, if the initial cloud consists of two spin-aligned states $\varphi_{211}$, we have
@@
where the eigenstate wave functions are normalized to 1. Here we discuss only the case with two identical gravitational atoms. If the clouds surrounding the two black holes are in different atomic eigenstates or with nonparallel spin alignment, the initial condition can also be constructed in a similar way for equal-mass binaries. Going beyond equal-mass binaries is, however, nontrivial. The initial conditions for $c_{i}$ in Eq. (14) cannot be obtained analytically. Nevertheless, it can be obtained by projecting the initial wave function on the molecular eigenstates numerically. The initial conditions for clouds in unequal-mass binaries will be left for future investigation.
@@ V.2 Cloud depletion
Before forming a common envelope, clouds in binaries may undergo mass transfer. For example, a cloud that is initially in $\varphi_{1,{\rm 211}}$ starts to oscillate between $\varphi_{1,{\rm 211}}$ and $\varphi_{2,{\rm 21-1}}$ when the wave functions of the two states develop a noticeable overlap as the two black holes get closer. As a result, the cloud loses mass to the second black hole via $\varphi_{2,{\rm 21-1}}$. It is claimed in Ref. [49] that, for equal-mass binaries, the cloud depletes rapidly around $a\sim 60r_{B}$ due to mass transfer to the $\varphi_{{\rm 21-1}}$ mode of the companion black hole. In this case, the cloud cannot survive to form a common envelope. In the following, we will show that the depletion is not efficient, except for very large $q$.
@@
During a nonresonant evolution, we have
@@
where $\eta$ is the coupling strength, $\omega\approx\Delta\epsilon+2\Omega$ with $\Delta\epsilon$ being the energy split between $\varphi_{1,{\rm 211}}$ and $\varphi_{2,{\rm 21-1}}$, and the integral over $t^{\prime}$ takes into account the slowly changing of $\omega$ and $\eta$. For $q\neq 1$, we have $|\Delta\epsilon|\approx(1-q^{2})\alpha^{3}/8\gg|2\Omega|$, and the clouds deplete significantly only if
$\Gamma_{2,{\rm 21-1}}\frac{\eta^{2}}{\omega^{2}}\frac{\Omega}{\dot{\Omega}}\approx\frac{40}{3}\frac{(1+q)^{1/3}q^{8}}{(1-q^{2})^{2}}\alpha^{2}{\rm I}_{4}(a)\gtrsim 1,\,$ (37)
where we have averaged $|c_{2}|^{2}$ over a timescale much longer than $\sqrt{\omega^{2}+\eta^{2}}$ but shorter than $\Omega/\dot{\Omega}$. We also defined ${\rm I}_{4}(a)=\left(\frac{a}{r_{B}}\right)^{4}G^{2}M^{2}\eta^{2}/\alpha^{6}$ which depends only on $q$. Figure  5 shows the numerical value of ${\rm I}_{4}(a)$ for different $q$, from which we conclude that the clouds barely deplete unless $q$ is relatively large. For example, by requiring the maximum of the left-hand side of the inequality (37) to be less than $1$, we get $q>3.27$ for $\alpha=0.1$. The case of $q=1$ is especially, in the sense that $|\Delta\epsilon|\simeq 0$, and $\omega$ is dominated by $2\Omega$. [^4] To be precise, $|\Delta\epsilon|\propto\mu\alpha^{5}$ and depends on the black hole spin. In this case, the depletion condition becomes
@@
where ${\rm I}_{7}(a)=\left(\frac{a}{r_{B}}\right)^{7}G^{2}M^{2}\eta^{2}/\alpha^{6}$. As shown in Fig. 5, even for $q=1$, the clouds cannot completely deplete. Compared to the estimation in Ref. [49], $\omega$ differs by $2\Omega$, which is due to binary rotation that is not considered in Ref. [49]. Moreover, the mass transfer can be resonant at certain orbits as discussed in Ref. [33], which, however, do not deplete the cloud completely either. Therefore, we conclude that, for the general mass ratio, the clouds generically survive to form common envelopes in the late inspiral.
Figure 5: Integrals in Eqs.  (37) and  (38). The vertical axis shows ${\rm I}_{x}(a)=\left(\frac{a}{r_{B}}\right)^{x}G^{2}M^{2}\eta^{2}/\alpha^{6}$, where colors of curves represent conditions for various mass ratios ($q=0.5$, $0.75$, 11, $1.25$ and $1.5$ ) with corresponding powers.
@@ VI Implications on orbital evolution
$E_{\rm c}=M_{\rm c}\alpha^{2}\left[\sum_{i}|c_{i}|^{2}\tilde{\epsilon_{i}}+(1+q)^{1/2}\tilde{a}^{-3/2}\sum_{i,j}c_{i}^{*}c_{j}\tilde{\eta}_{ij}\right],$ (39)
where $M_{c}$ is the cloud mass, and $\tilde{\epsilon}_{i}$ and $\tilde{\eta}$ are the dimensionless molecular eigenenergy and coupling strength, for example, shown in Fig. 2 and Fig. 3, respectively. The cloud also contributes to the total angular momentum with
@@
Figure 6: Orbital evolution in the phase space. The upper plot shows the evolution trajectories in the absence of clouds, where the eccentricity always decreases during the orbital decay. The middle and lower plots show the evolution trajectories in the presence of the cloud, where we assume a cloud of $M_{c}=0.1M$ and $0.01M$, respectively, originating from two spin-aligned superradiance states $\varphi_{1,211}$ and $\varphi_{2,211}$, and undergoing adiabatic evolution. In these cases, the eccentricity may increase as the orbits decay.
@@
According to Eq. (42), the cloud plays an important role in the eccentricity evolution during orbital decay. In particular, the right-hand side of Eq. (42) can be positive in the presence of the cloud, resulting in increased orbital eccentricity. For example, we consider the cloud that initially consists of two spin-aligned atomic states $\varphi_{1,211}$ and $\varphi_{2,211}$. During adiabatic evolution, we expect that the coefficients $c_{i}$ are constant in time and that the orbital evolution can be obtained by solving Eqs. (41) and  (42). In Fig. 6, we plot the evolution trajectories, assuming $\tilde{M}_{c}=0.1$. We find that eccentricity may increase when $4<\tilde{a}<6$ and $10<\tilde{a}$.
Figure 7: Eccentricity pumping during molecular level transitions. The plots show orbital evolution near the molecular level transition from $\bar{\psi}_{211}$ to $\bar{\psi}_{210}$, assuming different cloud masses. The upper plot assumes an initial eccentricity of $0.2$ at $\tilde{a}=3$, while the lower plot assumes an initial eccentricity of $0.1$ at $\tilde{a}=3$. The gray dashed line located at $a\sim 2.79r_{B}$ represents the resonant radius of the transition. The evolution is obtained by solving Eqs. (41),  (42), and  (43) numerically.
The growing eccentricity may also occur during nonadiabatic evolution, where $c_{i}$ change significantly due to resonant transitions of the cloud. For the cloud considered above, there is a resonant transition from $\bar{\psi}_{211}$ to $\bar{\psi}_{210}$ at $\tilde{a}\simeq 2.79$. In Fig. 7, we show the eccentricity evolution near the resonant orbit by solving Eqs. (41),  (42), and  (43) numerically. We find that for $\tilde{M}_{c}>10^{-2}$, the eccentricity varies significantly during the transition, leaving an eccentricity of ${\cal O}(0.1)$ at $a\sim r_{B}$.
Assuming equal-mass binaries, the frequency of the gravitational wave emitted at $a\sim r_{B}$ can be estimated as
@@
Therefore, for stellar mass binaries, the superradiant cloud can give rise to a nonzero eccentricity detectable in the LIGO-Virgo-KAGRA (LVK) band. Interestingly, recent analysis of the LVK event GW200105 indicates a median orbital eccentricity of $e\sim 0.145$ at an orbital period of $0.1$s [61]. Our work provides a potential origin of the median eccentricity. [^5] Note that, in practice, the mass of the boson field is a constant, and $\alpha$ varies with black hole masses. In other words, one has $f_{B}\propto\mu^{3}M^{2}$. Nevertheless, we do not expect to have eccentricity excitation in binaries with heavier black holes, because the superradiant growth is suppressed or the cloud might have already depleted. Therefore, the eccentricity excitation mechanism could explain the observed eccentricity in GW200105, while being in consistence with other LVK events.
Note that the eccentricity excitation mechanisms discussed above are different from the one in Refs. [47, 48], which is caused by the level transitions between atomic states and happens only in an inclined orbit and typically at a relatively large orbital separation. As a result, the induced eccentricity hardly survives to the LVK band and cannot be detected with the current ground-based gravitational wave detectors.
@@ VII Conclusion and Discussion
In this work, we investigate the evolution of the cloud in the stage where the orbital separation becomes comparable to the cloud size. We describe the cloud evolution with the gravitational molecular eigenstates. With this framework, we discuss the common envelope formation of the superradiant clouds and their late evolution in binaries. We find that most of the time the cloud evolves adiabatically; i.e., each eigenstate slowly adjusts itself as the orbit decays. Resonant transitions between different eigenstates may also happen when the energy levels cross or if the orbital frequency matches the energy difference between the eigenstates.
We have also investigated the potential effects on the orbital evolution of the clouds. We find that for elliptical orbits, the eccentricity may increase when the orbital separation is greater than 5 times the Bohr radius. This result proposes a potential origin for the median eccentricity at a small orbital separation that was recently implied by the LVK event GW200105 [61]. The cloud-induced eccentricity could be further tested with future gravitational-wave observation, providing an alternative way of searching for the ultralight boson field with gravitational wave observations.
This work provides a framework for investigating cloud evolution and its effects on binary black hole mergers as environments. While in this work we focus on the evolution of bound states, the bound states could also couple to unbounded states through the perturbation $H^{\prime}$, thereby exciting bosonic radiation. This leads to cloud ionization and extra dissipation of the binary binding energy, which can also be treated within the present framework. Cloud ionization in comparable mass black hole binaries has recently been observed in numerical simulations and analyzed in Ref. [62]. In particular, Ref. [62] shows that ionization occur in eccentric orbits. In principle, cloud ionization may influence orbital evolution and should eventually be taken into account in a complete treatment. However, ionization becomes efficient only when the orbital separation approaches the Bohr radius, i.e., $\tilde{a}\sim 1$. By contrast, the eccentricity growth identified in this work could occurs at much larger separations, e.g., $\tilde{a}>20$ (cf. Fig. 6), well before ionization becomes relevant. We therefore do not expect ionization to significantly modify our results in the large-separation regime. At small orbital separations, however, cloud ionization should be properly accounted for. It may deplete the cloud efficiently and quantitatively change the strength of resonant eccentricity pumping (cf. Fig. 7). A detailed study of ionization in gravitational molecules will be presented in future work. Together with ionization, the cloud environmental effects on gravitational waves from binary black hole mergers are promising directions for future study.
@@ ACKNOWLEDGMENTS
The authors thank Richard Brito, Yifan Chen, and Giovanni Maria Tomaselli for useful discussions. J. Z. is supported by the scientific research starting grants from the University of Chinese Academy of Sciences (Grant No. 118900M061), the Fundamental Research Funds for the Central Universities (Grants No. E2EG6602X2 and No. E2ET0209X2), and the National Natural Science Foundation of China (NSFC) under Grant No. 12147103.
DATA AVAILABILITY
The data that support the findings of this article are openly available  [63].